# Hong Sheng Tan

### v1

@@
[…]
Landau’s constant related to the classical zero free region of Riemann zeta function.
@@
(November 3, 2024)
Your Document Title
Hong Sheng Tan
MAM2309005@XMU.EDU.MY
(November 3, 2024)
Landau’s constant related to a zero free region of Riemann zeta function
Hong Sheng Tan
MAM2309005@XMU.EDU.MY
(November 3, 2024)
@@ 1 Introduction
$\zeta(s)=\frac{\pi^{\frac{s}{2}}}{\Gamma\left(\frac{s}{2}\right)}\left(-\frac{1}{s}-\frac{1}{1-s}+\int_{1}^{\infty}\sum_{n=-1}^{\infty}e^{-\pi n^{2}x}\left(x^{s/2-1}+x^{(1-s)/2-1}\right)dx\right).$
@@
The zeros of the Riemann zeta function are critical in understanding the distribution of prime numbers. These zeros are classified into trivial zeros, occurring at the negative even integers $s=-2,-4,-6,\ldots$, and nontrivial zeros, which lie within the critical strip $10<\sigma<$ 1. The Riemann Hypothesis conjectures that all nontrivial zeros have a real part ${}\sigma=\frac{1}{2}$. For further proofs and theoretical background on the Riemann zeta function, one can refer to Apostol [Apo76].
To study the behavior of nontrivial zeros of the Riemann zeta function, the zero-free region of the Riemann zeta function was first established by de la Vallée Poussin [dlVP00] in 1899. The zero-free region is an area in the critical strip where the Riemann zeta function does not vanish. The classical zero-free region has the form:
@@
counts the number of primes less than $x$. Landau [vL08] proved that for any $\rho>R$:
$\pi(x)=\int_{2}^{x}\frac{dy}{\log y}+O\left(x(\log x)^{-1/2}e^{-\sqrt{\log x/\rho}}\right).$
@@
1899 de la Vallée Poussin [dlVP00] 30.47
1962 Rosser, Schoenfeld [SRS62] 17.52
1970 Stechkin [Ste70] 9.65
1975 Rosser, Schoenfeld [RS75] 9.65
2002 Ford [For02] 8.46
2005 Kadiri [Kad05] 5.69
2022 Mossinghoff, Trudgian, Yang [MTY22] 5.56
@@
Landau [vL08] demonstrated that for any $f\in C_{n}$, one can take:
@@
The problems to find the value of $V$ were studied by Westphal [Wes38], Stechkin [Ste70], Rosser and Schoenfeld [SRS62, RS75], French [Fre66], Kondrat’ev [Kon77, AK90], Reztsov [Rez86], Arestov [AK90, Are92], Mossinghoff and Trudgian [MT14]. Most of the works are dedicated to estimating the upper and lower bounds of $V$. As of now, the best two-sided estimates for $V$ are:
@@
To determine the exact value of $V_{2}$, French [Fre66] substituted $x=\cos\phi$ and expressed every cosine polynomial in $C_{2}$ in the form:
@@
On the other hand, Arestov use the following properties of nonnegative trigonometric polynomials to evaluate the exact value of $V_{n}$ for $n=3,4,5,6$. Firstly, he noticed that for any trigonometric polynomial $f_{n}\in C_{n}$, the Fejér inequality [PS78] holds:
@@
Arestov [Are92] used some inequalities of nonnegative trigonometric polynomials to find the exact value of $\chi_{n}(a)$ on some specific intervals. This enabled him to calculate the exact value of $V_{n}$ for $n=3,4,5,6.$ The values of $V_{n}$ for $n=2,3,4,5,6$ are summarized in Table 2.
@@
2 French [Fre66] 53.1390720
3 Arestov [Are92] 36.9199911
4 Arestov [Are92] 34.8992259
5 Arestov [Are92] 34.8992259
6 Arestov [Are92] 34.8992259
Table 2: Values of $V_{n}$ as determined by Arestov.
@@
Sections 3 to Section 8 will be devoted to finding the exact values of $V_{n}$ for $n$ from 4 to 8. By applying the Fejér-Riesz theorem [Fej16], the problem of finding $\chi_{n}(a)$ is transformed into an optimization problem in $\mathbb{R}^{n}$ and then solved using optimization techniques. The techniques and general procedure will be described in Section 3, while the key constraints and result of the optimization problems will be summarized in Section 4 to Section 8. In the last two sections, the following theorems will be presented.
@@ 2 The exact value of $V_{2}$ and $V_{3}$
2.1 Some properties of function $v(f)$.
@@
To establish this result, we first require some preliminary lemmas. Three of the following lemmas provide simple inequalities that will be used to compare the values of $v(f)$ and $v(\bar{f})$. Readers may choose to skip ahead to Theorem 2.12 and return to these lemmas when the inequalities are needed. For a smoother reading experience, the proofs of Lemma 2.9, Lemma 2.10 and Lemma 2.11 are presented in the Appendix A.
@@ 3 Methodology
The Fejer-Riesz theorem was first employed by Kondrat’ev [Kon77] to find the upper bound of $V_{8}$ in his paper. This method was later employed by Mossinghoff and Trudgian [MT14] to determine the upper bound of $V_{n}$ for $n=8,12,16,\ldots,40$. The Fejer-Riesz theorem translates the problem of finding the exact value of $V_{n}$ into a minimization problem on $\mathbb{R}^{n}$, subject to certain conditions. The statement of the Fejer-Riesz theorem is presented as follows:
@@
[Fej16] Let $g(\phi)=\sum_{k=0}^{n}b_{k}\cos k\phi$ be a trigonometric polynomial with real coefficients, which is nonnegative for every real $\phi$ and $b_{n}\neq 0$. Then, there exist real numbers $\alpha_{0},\alpha_{1},\alpha_{2},\ldots,\alpha_{n}$ such that
$g(\phi)=\left|\sum_{k=0}^{n}\alpha_{k}e^{ik\phi}\right|^{2}.$
@@
[Mun00] Every continuous function on a compact set attains its extreme values on that set.
@@ 3.1 Karush-Kuhn-Tucker Method
[BSS06] The Problem $P$ is said to satisfy Slater’s condition if we can find a vector $\mathbf{x}\in\mathbb{R}^{n}$ such that
@@
[BSS06] Let $f:R^{n}\rightarrow R$, and $g_{i}:R^{n}\rightarrow R$ for $i=$ $1,\ldots,m$. Consider Problem ${P}$, which seeks to minimize $f({\mathbf{x}})$ subject to $g_{i}({\mathbf{x}})\leq 0$ for $i=1,\ldots,j$ and $h_{i}(\mathbf{x})=0$ for $i=1,\ldots,k$. Suppose $\overline{\mathbf{x}}$ is a solution to this problem, and that $f,g_{i}$ and $h_{i}$ are all differentiable at $\overline{\mathbf{x}}$. Furthermore, assume the gradients of the constraints, $\nabla g_{i}(\overline{\mathbf{x}})$, are linearly independent for all $i$. If $\overline{\mathbf{x}}$ solves Problem $\mathrm{P}$ locally, there exist scalars $u_{i}$ for $i\in I$ such that:
@@ 3.2 Penalty Function
[BSS06] Consider the following problem:
@@ 3.3 General Procedure
To prove this, we revisit the approach used by Arestov and Kondratev [AK90], which provides a set of key inequalities that give $\chi_{n}(a)$ lower bounds. The following theorem gives a simple lower bound for $\chi_{n}(a)$.
@@
[AK90] For ${}a\in\left(1,2\cos\frac{\pi}{n+2}\right]$, the functions $\chi_{n}(a)$ for $n\geq 2$ possess the following property
@@
By choosing different nonnegativefunction $m(\phi)$ such that $M=1$, one could obtain several lower bounds of $\chi_{n}(a).$ More specifically, Arestov and Kondrat’ev [AK90] have employed the following functions
@@
[AK90] For ${}a\in\left(1,2\cos\frac{\pi}{n+2}\right]$, the functions $\chi_{n}(a)$ for $n\geq 2$ possess the following properties:
@@
Then, a Matlab-based algorithm is employed to compute the ${}\frac{p_{2}(a)}{(\sqrt{a}-1)^{2}}$. The algorithm utilizes the Newton-Raphson method in conjunction with a penalty function approach to iteratively find the minimizer of the penalized objective function. The detailed Matlab implementation is provided in Appendix B. The Figure 5 shows the ratio ${}\frac{p_{2}(a)}{(\sqrt{a}-1)^{2}}$.
@@
Then, a Matlab-based algorithm is employed to compute the ${}\frac{q_{2,i}(a)}{(\sqrt{a}-1)^{2}}$ for $i=1,2$. The algorithm utilizes the Newton-Raphson method in conjunction with a penalty function approach to iteratively find the minimizer of the penalized objective function. The detailed Matlab implementation is provided in Appendix B.
@@
Then, a Matlab-based algorithm is employed to compute the ${}\frac{r_{2,i}(a)}{(\sqrt{a}-1)^{2}}$ for $i=1,2,3,4$. The algorithm employs the Newton-Raphson method, combined with a penalty function approach to iteratively minimize the penalized objective function (see Appendix B for the full Matlab implementation).
@@
A Matlab-based algorithm is subsequently used to calculate ${}\frac{s_{2,i}(a)}{(\sqrt{a}-1)^{2}}$ for $i=1,2,3,4$, by adopting the Newton-Raphson method in conjunction with a penalty function technique to iteratively find the minimizer of the penalized objective function (see Appendix B for the detailed Matlab implementation).

### v2

# Optimal Cosine Polynomials for Riemann Zeta Zero-Free Region

@@
Optimal Cosine Polynomials for Riemann Zeta Zero-Free Region
@@
(February 6, 2025)
@@ 1 Introduction
$\zeta(s)=\frac{\pi^{\frac{s}{2}}}{\Gamma\left(\frac{s}{2}\right)}\left(-\frac{1}{s}-\frac{1}{1-s}+\int_{1}^{\infty}\sum_{n=-1}^{\infty}\exp{(-\pi n^{2}x)}\left(x^{s/2-1}+x^{(1-s)/2-1}\right)dx\right),$
@@
The zeros of the Riemann zeta function are critical in understanding the distribution of prime numbers. These zeros are classified into trivial zeros, occurring at the negative even integers $s=-2,-4,-6,\ldots$, and nontrivial zeros, which lie within the critical strip $10<\sigma<$ 1. The Riemann Hypothesis conjectures that all nontrivial zeros have a real part ${}\sigma=\frac{1}{2}$. For further proofs and theoretical background on the Riemann zeta function, one can refer to Apostol [1].
To study the behavior of nontrivial zeros of the Riemann zeta function, the zero-free region of the Riemann zeta function was first established by de la Vallée Poussin [5] in 1899. The zero-free region is an area in the critical strip where the Riemann zeta function does not vanish. The classical zero-free region has the form:
@@
counts the number of primes less than $x$. Landau [10] proved that for any $\rho>R$:
$\pi(x)=\int_{2}^{x}\frac{dy}{\log y}+O\left(x(\log x)^{-1/2}\exp{\left(-\sqrt{\log x/\rho}\right)}\right).$
@@
1899 de la Vallée Poussin [5] 30.47
1962 Rosser, Schoenfeld [16] 17.52
1970 Stechkin [18] 9.65
1975 Rosser, Schoenfeld [17] 9.65
2005 Kadiri [8] 5.69
2022 Mossinghoff, Trudgian, Yang [12] 5.56
@@
Landau [10] demonstrated that for any $f\in C_{n}$, one can take:
@@
The problems to find the value of $V$ were studied by Westphal [19], Stechkin [18], Rosser and Schoenfeld [16, 17], French [7], Kondrat’ev [9, 3], Reztsov [15], Arestov [3, 2], Mossinghoff and Trudgian [11]. Most of the works are dedicated to estimating the upper and lower bounds of $V$. As of now, the best two-sided estimates for $V$ are:
@@
To determine the exact value of $V_{2}$, French [7] substituted $x=\cos\phi$ and expressed every cosine polynomial in $C_{2}$ in the form:
@@
On the other hand, Arestov use the following properties of nonnegative trigonometric polynomials to evaluate the exact value of $V_{n}$ for $n=3,4,5,6$. Firstly, he noticed that for any trigonometric polynomial $f_{n}\in C_{n}$, the Fejér inequality [14] holds:
@@
Arestov [2] used some inequalities of nonnegative trigonometric polynomials to find the exact value of $\chi_{n}(a)$ on some specific intervals. This enabled him to calculate the exact value of $V_{n}$ for $n=3,4,5,6.$ The values of $V_{n}$ for $n=2,3,4,5,6$ are summarized in Table 2.
@@
2 French [7] 53.1390720
3 Arestov [2] 36.9199911
4 Arestov [2] 34.8992259
5 Arestov [2] 34.8992259
6 Arestov [2] 34.8992259
Table 2: Values of $V_{n}$ as determined by French and Arestov.
@@
Sections 3 to Section 8 will be devoted to finding the exact values of $V_{n}$ for $n$ from 4 to 8. By applying the Fejér-Riesz theorem [6], the problem of finding $\chi_{n}(a)$ is transformed into an optimization problem in $\mathbb{R}^{n}$ and then solved using optimization techniques. The techniques and general procedure will be described in Section 3, while the key constraints and result of the optimization problems will be summarized in Section 4 to Section 8. In the last two sections, the following theorems will be presented.
@@ 2 The exact value of $V_{2}$ and $V_{3}$
2.1 Some properties of linear functional $v(f)$.
@@
To establish this result, we first require some preliminary lemmas. Three of the following lemmas provide simple inequalities that will be used to compare the values of $v(f)$ and $v(\bar{f})$. Readers may choose to skip ahead to Theorem 2.12 and return to these lemmas when the inequalities are needed. For a smoother reading experience, the proofs of Lemma 2.9, Lemma 2.10 and Lemma 2.11 are presented in the A.
@@ 3 Methodology
The Fejer-Riesz theorem was first employed by Kondrat’ev [9] to find the upper bound of $V_{8}$ in his paper. This method was later employed by Mossinghoff and Trudgian [11] to determine the upper bound of $V_{n}$ for $n=8,12,16,\ldots,40$. The Fejer-Riesz theorem translates the problem of finding the exact value of $V_{n}$ into a minimization problem on $\mathbb{R}^{n}$, subject to certain conditions. The statement of the Fejer-Riesz theorem is presented as follows:
@@
[6] Let $g(\phi)=\sum_{k=0}^{n}b_{k}\cos k\phi$ be a trigonometric polynomial with real coefficients, which is nonnegative for every real $\phi$ and $b_{n}\neq 0$. Then, there exist real numbers $\alpha_{0},\alpha_{1},\alpha_{2},\ldots,\alpha_{n}$ such that
$g(\phi)=\left|\sum_{k=0}^{n}\alpha_{k}\exp{(ik\phi)}\right|^{2}.$
@@
[13] Every continuous function on a compact set attains its extreme values on that set.
@@ 3.1 Karush-Kuhn-Tucker Method
[4] The Problem $P$ is said to satisfy Slater’s condition if we can find a vector $\mathbf{x}\in\mathbb{R}^{n}$ such that
@@
[4] Let $f:R^{n}\rightarrow R$, and $g_{i}:R^{n}\rightarrow R$ for $i=$ $1,\ldots,m$. Consider Problem ${P}$, which seeks to minimize $f({\mathbf{x}})$ subject to $g_{i}({\mathbf{x}})\leq 0$ for $i=1,\ldots,j$ and $h_{i}(\mathbf{x})=0$ for $i=1,\ldots,k$. Suppose $\overline{\mathbf{x}}$ is a solution to this problem, and that $f,g_{i}$ and $h_{i}$ are all differentiable at $\overline{\mathbf{x}}$. Furthermore, assume the gradients of the constraints, $\nabla g_{i}(\overline{\mathbf{x}})$, are linearly independent for all $i$. If $\overline{\mathbf{x}}$ solves Problem $\mathrm{P}$ locally, there exist scalars $u_{i}$ for $i\in I$ such that:
@@ 3.2 Penalty Function
[4] Consider the following problem:
@@ 3.3 General Procedure
To prove this, we revisit the approach used by Arestov and Kondratev [3], which provides a set of key inequalities that give $\chi_{n}(a)$ lower bounds. The following theorem gives a simple lower bound for $\chi_{n}(a)$.
@@
[3] For ${}a\in\left(1,2\cos\frac{\pi}{n+2}\right]$, the functions $\chi_{n}(a)$ for $n\geq 2$ possess the following property
@@
By choosing different nonnegativefunction $m(\phi)$ such that $M=1$, one could obtain several lower bounds of $\chi_{n}(a).$ More specifically, Arestov and Kondrat’ev [3] have employed the following functions
@@
[3] For ${}a\in\left(1,2\cos\frac{\pi}{n+2}\right]$, the functions $\chi_{n}(a)$ for $n\geq 2$ possess the following properties:
@@
Then, a Matlab-based algorithm is employed to compute the ${}\frac{p_{2}(a)}{(\sqrt{a}-1)^{2}}$. The algorithm utilizes the Newton-Raphson method in conjunction with a penalty function approach to iteratively find the minimizer of the penalized objective function. The detailed Matlab implementation is provided in B. The Figure 5 shows the ratio ${}\frac{p_{2}(a)}{(\sqrt{a}-1)^{2}}$.
@@
Then, a Matlab-based algorithm is employed to compute the ${}\frac{q_{2,i}(a)}{(\sqrt{a}-1)^{2}}$ for $i=1,2$. The algorithm utilizes the Newton-Raphson method in conjunction with a penalty function approach to iteratively find the minimizer of the penalized objective function. The detailed Matlab implementation is provided in B.
@@
Then, a Matlab-based algorithm is employed to compute the ${}\frac{r_{2,i}(a)}{(\sqrt{a}-1)^{2}}$ for $i=1,2,3,4$. The algorithm employs the Newton-Raphson method, combined with a penalty function approach to iteratively minimize the penalized objective function (see B for the full Matlab implementation).
@@
A Matlab-based algorithm is subsequently used to calculate ${}\frac{s_{2,i}(a)}{(\sqrt{a}-1)^{2}}$ for $i=1,2,3,4$, by adopting the Newton-Raphson method in conjunction with a penalty function technique to iteratively find the minimizer of the penalized objective function (see B for the detailed Matlab implementation).